The ratio of two numbers is 1:4, and the sum of these numbers is 40. Find the numbers.
Answer by richwmiller(17219)

x=1/4y
x+y=40
substitute
y+1/4y=40
multiply by 4
4y+y=160
5y=160
y=160/5=320/10=32
1/4*32=8=x
check
32+8=40
ok